### v1

# Machine-learned models for magnetic materials

Abstract
We present a general framework for modeling materials using deep neural networks. Material represented by multidimensional characteristics (that mimic measurements) is used to train the neural autoencoder model in an unsupervised manner. The encoder is trying to predict the material parameters of a theoretical model, which is then used in a decoder part. The decoder, using the predicted parameters, reconstructs the input characteristics. The neural model is trained to capture a synthetically generated set of characteristics that can cover a broad range of material behaviors, leading to a model that can generalize on the underlying physics rather than just optimize the model parameters for a single measurement. After setting up the model we prove its usefulness in the complex problem of modeling magnetic materials in the frequency and current (out-of-linear range) domains simultaneously.
Index Terms:
materials modeling, deep neural networks, synthetic data
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Modeling different phenomena or materials is always connected with a need to tune model-specific parameters. After developing a theory that describes the underlying physics, we come to the stage where it is necessary to choose the parameters of the model, most often adjusting them to obtain a correspondence between the model characteristics and the measurements. Depending on the model complexity, the typical approach is to optimize parameters by curve fitting using least-squares or one of the vast variety of optimization algorithms, such as gradient-based methods[1], simulated annealing[2], particle swarm optimizations[3, 4], or evolutionary algorithms[5, 6]. Regardless of the algorithm chosen, optimization is always the selection of a parameter for a given set of measurement data.
With the advent of big data, deep learning, i.e. paradigm that uses deep neural networks (NNs), revolutionary breakthroughs in many conventional machine learning and pattern recognition tasks are a fact, such as computer vision or signal processing [7, 8]. The accompanied data-driven materials modeling approach [9, 10, 11, 12] also provides significant shift in materials science methodology from simple trial-and-error routine to intelligent discovery of new materials with demanding properties [13, 14, 15, 16, 17].
In this paper, we propose a similar approach but for learning any analytical model of a material in terms of intelligent fitting of its parameters. In the following, we train a machine learning model composed of a deep NN to be able to select parameters of a given theoretical model over a wide range of material characteristics at the same time, not just to single measurement data, which from machine learning perspective would be just overfilling of this single characteristic. Our goal is to build a model capable of generalizing over a variety of characteristics than a single one, which also means that the NN is able to capture the physics behind the theoretical model.
II Model and data generation
To show the capability of the proposed approach, we will focus on a simple but very useful phenomenological model of magnetic materials in a wide frequency range capturing linear but also nonlinear regimes [18]. Various approaches to modeling magnetic materials in nonlinear conditions were proposed, including famous Jiles-Atherton (JA) model [19, 20], models based on Finite Element Method (FEM) that can describe local spatial dependence of magnetization [21, 22], or lumped element equivalent circuits (LEEC) suitable for direct implementation in SPICE or EMTP simulators [23, 24].
Here, we will use LEEC representing impedance characteristics of magnetic rings in both frequency and current domains, introduced in [18], and further analyzed in terms of its parameters optimization [25], also using AI methods [26]. Nanocrystalline rings are currently used in many power and electronic devices [27] to meet the requirements related to electromagnetic compatibility by suppressing electromagnetic interferences occurring in these devices and the requirements of insulation coordination by suppressing overvoltages occurring in these devices, in particular, overvoltages occurring in Gas-Insulated Switchgear. The problem of overvoltage suppression in GIS is important in the case of high-voltage GIS starting from 550 kV and ultra-high voltage GIS up to 1100 kV [28], in which fast-changing overvoltages (with frequencies of several MHz) overlap with slow-changing waveforms (with frequencies of 50/60 Hz), which are approximately constant. The significance of these conditions comes from the fact that the peak values of these overvoltages determine the design of the insulation systems of the GIS devices.
To make our analysis realistic, we collect a set of measurements that will serve as a test set for the developed method. In Fig. 1(a), there is presented a family of impedances as a function of frequency $f$, taken for different bias current $I_{\mathrm{DC}}$ which controls the ring magnetization: from a linear range up to the ring saturation. The color of the measured impedance curves $Z^{\mathrm{meas}}$, that changes gradually from red, for $I_{\mathrm{DC}}=0$ A, to blue, for $I_{\mathrm{DC}}=27$ A, ($N_{\mathrm{bias}}=31$ different values in total) shows impedances for different biasings, combining this way multiple frequency characteristics controlled by the bias current $I_{\mathrm{DC}}$: $Z^{\mathrm{meas}}=Z^{\mathrm{meas}}(I_{\mathrm{DC}},f)$. Having families of characteristics, not just a single one, will be crucial when proving the model’s capability to capture broad physics rather than just fitting a single feature. Methods for measuring magnetic rings over a wide range of frequencies and saturations are not a trivial task (especially reaching the saturation limit is challenging) and are described in detail in Ref. [29].
Figure 1: Impedance of a magnetic ring (a) measured (inset) as a function of frequency for various biasings that controls operating point from linear range ($I_{\mathrm{DC}}=0$ A, red curve) to the full saturation ($I_{\mathrm{DC}}=27$ A, blue curve), and (b) generated using lumped element model (inset) with continuous transition between different curves (operating regimes) and randomly selected outermost characteristics (red and blue one). The generated data will serve as a training set for the introduced neural network model, while the measurement ones will be used to test the already trained NN model.
Let us now introduce the analytical model of a ring in the form of LEEC impedance synthesized out of the series of $M=7$ LR elements connected in parallel, as presented in Fig. 1(b), with the equivalent impedance in a form
$Z^{\mathrm{model}}(s)=\sum_{i=1}^{M}\frac{sR_{i}}{s+\frac{R_{i}}{L_{i}}},$ (1)
where $s=j2\pi{}f$, and $j$ is the imaginary unit. The goal for the NN will be to learn the model defined in Eq. 1 and predict its parameters, i.e., to find $\{L_{i},R_{i}\}$ for a given frequency characteristics over the wide range of their possible behaviors. Therefore, to train the NN model and to force it to generalize over different behaviors, we have to synthesize the training dataset using the analytical model with random parameters. Generating data and training neural models on artificially synthesized datasets with controlled properties is currently a growing trend in machine learning [30].
[…]
III Neural network model
After defining the generated training and measured test datasets, let us define the NN responsible for finding the model parameters. The network will be composed of two parts: encoder and decoder – depicted in Fig. 2, and trained in an unsupervised manner. The role of the encoder is to predict the analytical model parameters $\{L_{i},R_{i}\}$ (their values define the latent space of our autoencoder), while the decoder will just calculate the frequency characteristic directly using Eq. 1 with parameters (latent space) found by the encoder.
Figure 2: Neural network basic model with an encoder-decoder structure, which is of fundamental type and used in various fields of machine learning or image processing. The encoder is made of three convolutional layers (CNN) and three fully connected (FC) layers. The CNN layers are used as feature extractors, while the FC layers are used for parameters ($\{L_{i},R_{i}\}$) predicting based on previously found representations. The decoder ($Z$) just implements an analytical formula for the impedance of the LEEC model. The goal of the network is to reconstruct the input characteristics at the output.
The goal for the NN is to reconstruct input characteristics as best as possible, with the smallest reconstruction error called loss – this means that in our training, we do not need any labels and only frequency characteristics to be reconstructed which makes our training scheme fully unsupervised.
The standard solution would also be to use an NN as the decoder, but this approach makes less sense because we need the decoder only to train the encoder, which in turn, is essential for us (it predicts the analytical model parameters). Therefore, we successfully replaced the NN decoder with the analytical model itself (here defined by Eq. 1 for equivalent LEEC impedance) implemented as part of the NN (without trainable parameters, but using the model parameters ($\{L_{i},R_{i}\}$) predicted by the encoder) to be able to use the standard backpropagation of loss mechanism during the training.
Frequency characteristics are organized on a 1D evenly spaced grid along multiple decades on the log scale ($50$ points from $135$ Hz to $10^{7}$ Hz). To build an effective neural representation, we used convolutional neural network (CNN) layers as feature extractor parts, which is a common way to process data organized on regular grids [31, 32]. Then, we applied several fully connected (FC) layers serving as parameter predictors, which are based on previously found representations (by CNN). The detailed structure of the encoder network is presented in Fig. 2: three CNN feature extractor layers are followed by three FC layers with $2M$ outputs in the last layer that predict the parameters. We select CNN as a feature extractor (backbone) part, which is common practice in signal or image processing [8, 7]; however, in principle, any other type of network could serve as backbone network at this stage, depending on applications – such as recurrent neural networks or nowadays very popular transformer networks [33] with a self-attention mechanism incorporated.
To train a NN model, we have to define a loss function $\mathscr{L}$, which in our case will be just the decoder (reconstruction) loss $\mathscr{L}=\mathscr{L}_{\mathrm{dec}}$, i.e. the mean squared difference between input and output impedance, taken in logarithmic scale for real and imaginary part separately:
$\mathscr{L}_{\mathrm{dec}}=\mathrm{MSLE}\left(\Re(Z),\Re(\hat{Z})\right)+\mathrm{MSLE}\left(\Im(Z),\Im(\hat{Z})\right),$ (2)
with $\hat{Z}$ meaning reconstructed characteristic (decoder output), and mean squared logarithmic error defined as
$\mathrm{MSLE}\left(Z,\hat{Z}\right)=\frac{1}{N}\sum_{j=1}^{N}\log^{2}\!\left(\frac{Z_{j}}{\hat{Z}_{j}}\right).$ (3)
The autoencoder is trained over the whole training set and the sum in Eq. 2 goes over $N=N_{\mathrm{train}}\times N_{\mathrm{bias}}$ frequency characteristics $Z$ in the training set. However, in practice, during a single training step of the stochastic gradient descent (SGD) method, only a smaller random portion (mini-batch) is used, here $N_{\mathrm{batch}}=1024$. Training typically lasted $100$-$500$ epochs up to the loss saturation. To validate the NN model performance after the training, we also define the relative error of the fitting (FE):
$\mathrm{FE}=\sqrt{\mathrm{RE}\!\left(\Re(Z),\Re(\hat{Z})\right)+\mathrm{RE}\!\left(\Im(Z),\Im(\hat{Z})\right)},$ (4)
with
$\mathrm{RE}\left(Z,\hat{Z}\right)=\frac{1}{N}\sum_{j=1}^{N}\left(\frac{\log(Z_{j})-\log(\hat{Z}_{j})}{\log(Z_{j})}\right)^{2}.$ (5)
IV Basic model results
Let us now analyze the NN model performance. To verify the model’s ability to generalize over the broad variety of characteristics we estimate the validation error using Eq. 5 but calculated after the training on some additional validation set generated in the same way as the training one. Table I proves that the model is capable of predicting parameters for a variety of characteristics (as example one presented in Fig. 1(b)), which is a crucial property, a necessary condition that must be met in order to have a chance of predicting parameters for unknown measured characteristic.
models validation error
basic Siamese modified Siamese
8.24% 5.29% 4.54%
models test error
basic Siamese modified Siamese
12.54% 7.22% 5.18%
Table I: Relative fitting error for different modifications of the NN model calculated using validation synthetic set and measured test set.
The result presented in 1st column of Table I shows that the basic version of the NN model (depicted in Fig. 2) generalizes quite well with the fitting error slightly above $8\%$.
Figure 3: Results for testing the basic neural network model on the measured ring data: (a,b) examples of fitting the analytical model (solid curves), with parameters predicted by the NN model, versus measurements (dots) for the selected bias currents $I_{\mathrm{DC}}=\{0.9,6.9,12.5\leavevmode\nobreak\ \mathrm{A}\}$, (c) relative fitting error for the various bias currents $I_{\mathrm{DC}}$ plotted versus frequency, and (d) characteristic frequencies of the predicted parameters (each representing LR pair in the LEEC series) versus bias current $I_{\mathrm{DC}}$.
Let us now perform tests and apply the model to realistic, measured data as presented in Fig. 1(a). Comparison of characteristics for the analytical model (defined in Eq. 1, with parameters predicted via the NN model) and the measured one, for selected $I_{\mathrm{DC}}$ values, are presented in Fig. 3(a,b) and shows quite good fitting performance over the broad range of bias currents. It is worth recalling here that the measurement data was not a part of the training dataset. Moreover, we calculated the relative fitting error systematically, using Eq. 5 but now on the measured data, for every bias current over the frequency range as presented in Fig. 3(c). In general, we observe that the error does not exceed $30\%$, which is quite a resonable result for fitting the model over so wide ranges of frequencies, impedance amplitudes, and bias currents $I_{\mathrm{DC}}$. On average it is even smaller and equals $12.54\%$.
V Siamese neural network model
The results obtained for the basic NN model are acceptable but have one substantial drawback. If we look at Fig. 3(d) which presents values of characteristic frequencies defined as $\omega_{i}=R_{i}/L_{i}$ for each of LR pairs in a series (of size $M=7$) for subsequent $I_{\mathrm{DC}}$ currents, we observe that the values are irregular and quite independent from each other which seems to be unphysical. In order to restore the more realistic behavior which is more likely to be continuous (we do not expect to have any phase transitions in the model here), the single NN model was extended to the Siamese model shown in Fig. 4.
Figure 4: Siamese neural network model composed of two copies of the basic model that share the weights and process two neighbor characteristics (that differ in bias current by $\Delta I$) at the same time enabling to train the model to predict similar (or continuously changing) parameters for similar characteristics.
The idea behind the Siamese NNs, popular in many ML areas, such as face recognition [34], is to have two (or more) copies of a network that works in parallel (but shares weights) and process different data samples at the same time. Processing multiple samples at once enables to regularize of the model behavior with respect to the similarity between data samples by adding an extra term to the loss function. In our case we can process neighbor characteristics, i.e. for $I_{\mathrm{DC}}$ and $I_{\mathrm{DC}}+\Delta I$ – see Fig. 4, and enforce parameters predicted for that data points to change continuously, by including additional loss terms in a form:
$l_{\mathrm{cont}}(L)=\frac{1}{N_{\mathrm{bias}}-1}\sum_{k=1}^{N_{\mathrm{bias}}-1}\log^{2}\!\left(\frac{L_{k+1}}{L_{k}}\right).$ (6)
The sum in Eq. 6 goes for every $N_{\mathrm{bias}}-1$ neighboring pair in a given characteristics family. It penalizes too high a change of the parameters between neighbors. The total continuity loss should include $M$ parameters and all $N_{\mathrm{train}}$ training families:
$\mathscr{L}_{\mathrm{cont}}(L)=\frac{1}{MN_{\mathrm{train}}}\sum_{i=1}^{M}\sum_{j=1}^{N_{\mathrm{train}}}l_{\mathrm{cont}}(L_{i,j}).$ (7)
In the same way, we define $\mathscr{L}_{\mathrm{cont}}(R)$. Moreover, we found that in our case it is beneficial to add yet another term to the loss function:
[…]
in order to maintain a monotonic decrease in the parameter values with the bias current $I_{\mathrm{DC}}$, which is justified by the ring material saturation. The total monotonicity loss $\mathscr{L}_{\mathrm{mono}}(L)$ is defined in the same way as in Eq. 7. All in all, the loss function for the Siamese NN model is as follows:
$\mathscr{L}_{\mathrm{Siam}}=\mathscr{L}_{\mathrm{dec}}+$ $\alpha_{1}\mathscr{L}_{\mathrm{cont}}(L)+\alpha_{2}\mathscr{L}_{\mathrm{cont}}(R)+$
$+$ $\alpha_{3}\mathscr{L}_{\mathrm{mono}}(L)+\alpha_{4}\mathscr{L}_{\mathrm{mono}}(R),$ (9)
with the training hyperparameters $\alpha_{i}$ that control the impact of the loss components responsible for keeping the continuity of the parameters under $I_{\mathrm{DC}}$ changes.
If we look at the results for the Siamese NN model presented in Fig. 5(a,b), where we force the model to predict similar parameters for similar characteristics, we can observe that now characteristic frequencies $\omega_{i}=L_{i}/R_{i}$ change with the bias current $I_{\mathrm{DC}}$ continuously and in some case monotonically – see Fig. 5(b). Moreover, a surprising but also very desirable property is that forcing the continuity for the change of the parameters $\{L_{i},R_{i}\}$ with $I_{\mathrm{DC}}$ at the same time has made fitting of the analytical model better than before – in Fig. 5(a) the relative fitting error typically does not exceed $15\%$ with the average value over frequencies and currents equals $7.22\%$.
Figure 5: Same as in Fig. 3 but results for (a,b) Siamese neural network, and (c,d) modified Siamese model.
VI Modified Siamese model
Fig. 5(b) presents yet another feature of the Siamese model (which was also partially present in the basic NN model) that it has a strong tendency to select features-parameters of the predicted analytical model (Eq. 1). It manifests in the fact that some groups of parameters, i.e. LR pairs in a series, are irrelevant – their characteristic frequencies, here $\omega_{1}$, $\omega_{2}$, $\omega_{3}$, lie outside the ring working frequencies, which means that they can be safely skipped in the model. This motivates us to reduce the dimensionality of the model by assuming $M=5$. Moreover, we observed that enforcing a more uniform positioning of the characteristic frequencies in the frequency domain over which the model operates can be beneficial in terms of further reducing the model size. Therefore, we added another component to the loss function of the form:
$\mathscr{L}_{\omega}=\frac{1}{NM}\sum_{j=1}^{N}\sum_{i=1}^{M}\left(\frac{R_{i}}{L_{i}}-\Omega_{i}\right)^{2},$ (10)
with $N=N_{\mathrm{train}}\times N_{\mathrm{bias}}$, $M=5$, and the constant frequency nodes centered at positions $\Omega_{i}=\{10^{3},10^{4.75},10^{6.5},10^{8.25},10^{10}\}$. Now the loss function for the modified Siamese NN model is the following:
$\mathscr{L}_{\mathrm{Siam^{\prime}}}=\mathscr{L}_{\mathrm{Siam}}+\alpha_{5}\mathscr{L}_{\mathrm{\omega}}.$ (11)
The results in Fig. 5(c,d) clearly show that now the positioning of the characteristic frequencies is more homogeneous and all of $\omega_{i}$ occupies ring working frequency range. In addition, the relative fitting error obtained here is less than $10\%$ in most cases and on average is the smallest among all three NN models, reaching $5.18\%$.
VII Discussion and Conclusions
In the first approach to the problem, we defined a network that predicts the parameters directly (containing the encoder part only) and compares them with the parameters used to generate a given characteristic. Such training was strictly supervised by using the encoder loss, defined as
$\mathscr{L}_{\mathrm{enc}}=\frac{1}{2NM}\sum_{j=1}^{N}\sum_{y_{i}\in\{L_{i},R_{i}\}}^{2M}\log^{2}\left(\frac{y_{ij}}{\hat{y}_{ij}}\right).$ (12)
The sum in Eq. 12 goes over the training samples and parameters: $\hat{y}_{ij}$ means the $i$-th parameter ($L_{i}$ or $R_{i}$) predicted for $j$-th sample, $Z_{j}$, whereas $y_{ij}$ represents ground truth parameters – we already know them from the generation process. However, this approach simply failed because it was not possible to effectively train a neural model that would generalize beyond a limited set of characteristics being some narrow part of the training data. Much better results came from the unsupervised training where the autoencoder model is simply given a variety of characteristics and has to attempt to reconstruct them. The model proposed in our scheme is able to generalize well, which is shown in Table I where the validation error reaches a few percent for the training set with a huge variety of characteristics. Consequently, the NN model is able to correctly fit parameters for previously unseen measured test data. At the same time, it is worth noting that the diversity of the training set fulfilled a kind of regularization, ensuring that the NN does not overfit a particular behavior.
In the proposed framework the materials are represented by multidimensional features, in our case as normalized impedance data in the form of real and imaginary parts calculated from the modulus and phase values of the data generated. The way the NN is trained is designed so that we can naturally enforce continuous (or any other physical) behavior on the entire family of measurement characteristics simultaneously.
We also observe the important effect that imposing continuity of the predicted parameters as a function of the bias current using the Siamese network, i.e., forcing the transitions to be physical, causes the fit to individual characteristics in the family to improve significantly.
The NN model has also a robust tendency to parameter selection. In the experiment we have shown that too large model leads to a fit in which some components did not carry a significant contribution, thus neglecting these components effectively reduces the found analytical model while keeping its error equally small.
VIII Methods
Data generation was performed with the usage of Pytorch library [35] on GPU, which significantly speedups the generation process from a few hours (when using CPU) to a dozen seconds (GPU). NN model (encoder and decoder part) was implemented using Pytorch with an automatic differentiation engine (autograd) used to track the gradient of the loss with respect to the model parameters during the backpropagation stage. The training of the single model on the GPU Nvidia A6000 took about 1 hour.
Measurements were taken for NANOPERM nanocrystalline rings from Magnetec GmbH labeled M-676 [36], for different magnetization currents $I_{\mathrm{DC}}$ up to ring saturation using the Keysight Technologies Impedance Analyzer E4990A-030 to measure the impedance frequency characteristic and Keysight Technologies DC Power Supply N8731A to supply the DC-bias current $I_{\mathrm{DC}}$.
IX Acknowledgments
The authors would like to acknowledge the financial support of Project No. [anonymized] granted by [anonymized].